OpenStudy (anonymous):

Simplify: fourteen times j plus 2 times j plus twenty-two times k minus 7 times j minus fourteen times k plus 3 times k 9 times j plus eleven times k 9 times j minus eleven times k thirty-one times j minus eleven times k thirty-one times j plus eleven times k

OpenStudy (sasogeek):

14j + 2j + 22k - 7j - 14k + 3k simplify this and tell me what u get

OpenStudy (anonymous):

look mhann jus give mhe da answer

OpenStudy (sasogeek):

why? what if i give u a wrong answer? will u walk away with it?

OpenStudy (sasogeek):

it's simple addition and subtraction... come on now... at least try it

OpenStudy (anonymous):

jus foget it brahh #gone dhipp

OpenStudy (sasogeek):

please don't tell me you're that lazy... ?

OpenStudy (sasogeek):

addition and subtraction..... seriously u can't do that? group like terms and add them

OpenStudy (anonymous):

plz dnt tell mhe u suk dhick

OpenStudy (sasogeek):

no i don't, i'm sorry dude, no one's gonna help u if you're not ready to learn ;) learn it up bro, just add the like terms right there and you're done ;)

OpenStudy (anonymous):

Hey Michael, you should try that yourself. If you keep practicing you won't feel lazy when you see another expression like that. Also, we are here to help you find the answer, not to hand it to you.

OpenStudy (anonymous):

BOTH YALL DHIPP 6IHH

OpenStudy (sasogeek):

I'll group the like terms for u, u do the addition and them tell me what the answer is

OpenStudy (anonymous):

UMMMHMMM

OpenStudy (sasogeek):

14j + 2j + 22k - 7j - 14k + 3k 14j+2j-7j+22k-14k+3k continue to find the answer

OpenStudy (anonymous):

BIHH SUK A FAT ONE

OpenStudy (sasogeek):

i didn't think you were that bad at maths, u need to learn, 14+2-7 =16-7=9 22-14+3=8+3=11 do u get what's going on?

OpenStudy (anonymous):

YEEE

OpenStudy (sasogeek):

if u don't learn this, i won't be there at ur test to help u and you're gonna fail if u don't learn, so do that now ;)

OpenStudy (anonymous):

YEA